# Influence of plume activity on thermal convection in a rectangular cell

### 13 Apr 2026 v1

@@ 3.1 Global transports of heat and momentum
The variation of $Nu$ with $Ra$ in the rectangular box is shown in figure 3(a). The best fit to the data in this range of $Ra$ follows $Nu=0.15Ra^{0.29}$. We also plot the Nusselt numbers from DNS studies in horizontally-periodic square cuboids of $\Gamma=32$ (Blass et al., 2021) and $\Gamma=4$ (van Reeuwijk et al., 2008), both for $Pr=1$, as well as data for $Pr=0.7$ in cylindrical cells of $\Gamma=1$  (Scheel et al., 2012; Wagner et al., 2012) and in a horizontally-periodic cuboid of $\Gamma=4$ (Samuel et al., 2024). $\Gamma=L/H$ is the aspect ratio. This power-law exponent is close to $2/7$ obtained in the mixing-zone model (Castaing et al., 1989) and to that by the turbulent boundary layer model (Shraiman & Siggia, 1990; Siggia, 1994). We conclude that the heat transport is largely insensitive to the aspect ratio (for $\Gamma\geq 0.5$) and the shape of the convection cell, and also to whether the domain is closed or horizontally periodic. The mixing model of Castaing et al. (1989) is linked to boundary layer only indirectly by being the source of plumes, whereas the Shraiman-Siggia model (Shraiman & Siggia, 1990) relies strongly on our empirical understanding of the turbulent boundary layer. The two models do not share much physics for a first look, yet they produce the same result on the NuRaNu-Ra relation. A plausible inference to draw from this observation is that the dynamical aspects of boundary layers and the mixing regions are intimately intertwined for these Rayleigh numbers.
@@ 3.3 Fluctuations in high and low activity bulk regions
We now compute the scalings of RMS temperature for the entire domain as
@@ 4.2 Boundary layer evolution on horizontal walls
We show in figures 9 and  10 the local boundary layer thicknesses computed from these profiles, normalised by that in the shear region, for normalised distances $x_{0}/L_{x}$. To avoid clutter, we show data only at the bottom plate. Figure 9 shows the variation of the VBL thickness estimated using the maximum method in panel (a), and that using the slope method in panel (b). Both methods yield local VBL thicknesses that increase, up to $x_{0}/L_{x}\approx 0.4$ in the wind direction, indicated using orange arrows. See also Qiu & Xia (1998a) and Wagner et al. (2012) for similar observations. Figure 9(b) shows that $\delta_{u}^{sl}(x_{0})$ starts to decrease as the ejection region is approached, but not the $\delta_{u}^{max}(x_{0})$.As the ejection region in the present configuration is far from the sidewalls, it is possible that the decline of the local VBL thickness is an artifact of the slope method.Clusters of thermal plumes with locally smaller spacing detach at the edge of the ejection region than in its center, where a stagnation region forms due to colliding mean flows. This behaviour can be probed only by the slope method that builds on vertical temperature derivatives at the wall. With increasing $Ra$, larger number of finer-scale plumes aggregates, causing broader and less coherent clustering regions, thus decreasing thickness ratios with increasing $Ra$.
@@
As shown in figure 10, TBL thicknesses from both methods increase all the way to $L_{x}/2$ up to $Ra=3\times 10^{6}$ but decreases with $x_{0}$ in the region $x_{0}\geq 0.4L_{x}$ for higher $Ra$. It is also clear that the variation of the TBL thickness over the plate becomes milder as $Ra$ increases, consistent with earlier observations (Lui & Xia, 1998; Pandey, 2021). Figure 9 further suggests that the VBL thickness varies over the plate more strongly than the TBL thickness.

### 04 May 2026 v2

@@ 3.1 Global transports of heat and momentum
The variation of $Nu$ with $Ra$ in the rectangular box is shown in figure 3(a). The best fit to the data in this range of $Ra$ follows $Nu=0.15Ra^{0.29}$. We also plot the Nusselt numbers from DNS studies in horizontally-periodic square cuboids of $\Gamma=32$ (Blass et al., 2021) and $\Gamma=4$ (van Reeuwijk et al., 2008), both for $Pr=1$, as well as data for $Pr=0.7$ in a horizontally-periodic cuboid of $\Gamma=4$ (Samuel et al., 2024). $\Gamma=L/H$ is the aspect ratio. This power-law exponent is close to $2/7$ obtained in the mixing-zone model (Castaing et al., 1989) and to that by the turbulent boundary layer model (Shraiman & Siggia, 1990; Siggia, 1994). We conclude that the heat transport is largely insensitive to the aspect ratio (for $\Gamma\geq 0.5$) and the shape of the convection cell, and also to whether the domain is closed or horizontally periodic. The mixing model of Castaing et al. (1989) is linked to boundary layer only indirectly by being the source of plumes, whereas the Shraiman-Siggia model (Shraiman & Siggia, 1990) relies strongly on our empirical understanding of the turbulent boundary layer. The two models do not share much physics for a first look, yet they produce the same result on the NuRaNu-Ra relation. A plausible inference to draw from this observation is that the dynamical aspects of boundary layers and the mixing regions are intimately intertwined for these Rayleigh numbers.
@@ 3.3 Fluctuations in high and low activity bulk regions
We now compute the RMS temperature for the entire domain as
@@ 4.2 Boundary layer evolution on horizontal walls
We show in figures 9 and  10 the local boundary layer thicknesses computed from these profiles, normalised by that in the shear region, for normalised distances $x_{0}/L_{x}$. To avoid clutter, we show data only at the bottom plate. Figure 9 shows the variation of the VBL thickness estimated using the maximum method in panel (a), and that using the slope method in panel (b). Both methods yield local VBL thicknesses that increase, up to $x_{0}/L_{x}\approx 0.4$ in the wind direction, indicated using orange arrows. See also Qiu & Xia (1998a) and Wagner et al. (2012) for similar observations. Figure 9(b) shows that $\delta_{u}^{sl}(x_{0})$ starts to decrease as the ejection region is approached, but not the $\delta_{u}^{max}(x_{0})$.As the ejection region in the present configuration is far from the sidewalls, it is possible that the decline of the local VBL thickness is an artifact of the slope method. Clusters of thermal plumes with locally smaller spacing detach at the edge of the ejection region rather than from its center, where a stagnation region forms due to colliding mean flows. This behaviour can be probed only by the slope method that builds on vertical temperature derivatives at the wall. With increasing $Ra$, larger number of finer-scale plumes aggregates, causing broader and less coherent clustering regions, thus decreasing thickness ratios with increasing $Ra$.
@@
As shown in figure 10, TBL thicknesses from both methods increase to $L_{x}/2$ up to $Ra=3\times 10^{6}$ but decreases with $x_{0}$ in the region $x_{0}\geq 0.4L_{x}$ for higher $Ra$. It is also clear that the variation of the TBL thickness over the plate becomes milder as $Ra$ increases, consistent with earlier observations (Lui & Xia, 1998; Pandey, 2021). Figure 9 further suggests that the VBL thickness varies over the plate more strongly than the TBL thickness.